OpenStudy (anonymous):

Help 3 ≤ |x| ≤ 4 I honestly have no idea how to begin. I need an interval notation set. Do I write the equations each having (-) sign on one number? then solve?

OpenStudy (dumbcow):

might help to look at it graphically 

OpenStudy (dumbcow):

algebraically, whenever you have absolute value split it into 2 cases \[|x| \ge 3\] \[\rightarrow x \ge 3 .... x \le -3\]

OpenStudy (anonymous):

I see. So I have to do it with 3 and 4...Yeah the drawing doesnt make sense to me at all.

OpenStudy (dumbcow):

haha oh well

OpenStudy (anonymous):

haha. sadly...anyway thanks

OpenStudy (anonymous):

 3 ≤ |x| ≤ 4 So... x≥3....x≤−3 X≤4....x≥-4 so... the answer would be [-4,-3] u [3,4]?